A string of length 1 m is fixed at one end with a Bob of mass 100 g and the string makes 2/pi rev /s around a vertical axis through a fixed point. The angle of inclination of the string with the vertical is 

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
 
w = 2pi*2/pi = 4 rad/s.
npw consider angle thethaa wth vertical and do comonents of Tension along horizontal and vertial and balnce them with mg and mw^2r
where from geomtery r = lsinthetha.
after balancing the eqns would be:
T = mw^2l
T = 100/1000 * 16 * 1 = 1.6 N
in vertical direcn.
Tcosthetha = mg
cosththa = 5/18
or
thetha = arccos(5/18)